OpenStudy (anonymous):

A car moving at 20 m/s slows down at -1.5 m/s^2 to a velocity of 10 m/s. How far did the car go during the slow down? Thank you so much.

OpenStudy (anonymous):

IS that meant to be delta

OpenStudy (anonymous):

\[20^{2}= 10^{2} + 2(-1.5)(distance)\]

OpenStudy (anonymous):

use \[{v_f}^2={v_0}^2+2ad\] \[\frac{400-100}{-3}=d=-100\]

OpenStudy (anonymous):

v f 2 =v 0 2 +2ad 400−100 −3 =d=−100 or 20 2 =10 2 +2(−1.5)(distance)

OpenStudy (anonymous):

Given, initial velocity, u = 20m/s velocity at particular time, v = 10m/s acceleration, a = -1.5 m/s^2 From equation of motion, \[v ^{2}-u ^{2}=2aS\]where S is the distance travelled therefore, \[(10)^{2}-(20)^{2}=2\times(-1.5)\times S\]100 - 400 = -3 S -300 = -3 S S=100 meter